mikey biked a distance of 75 miles in 6 hours. what speed was mikey traveling
Answer by stanbon(75887)

mikey biked a distance of 75 miles in 6 hours. what speed was mikey traveling
-----
rate = distance/time
-----
rate = 75/6 mph = 25/2 mph = 12.5 mph
-------------
Cheers,
Stan H.
-------------